OpenStudy (ray_sihota):

A random sample of students who rent apartments was taken. What is the probability of getting a sample mean monthly rent that exceeds $660 n=100 Slandered deviation=80 m= 650 x bar= 650 I did 660-650/80= 0.125 but the answer is 0.1056, what am i doing wrong

OpenStudy (faiqraees):

Since it is exceeds which is greater than you have to find the find the cumulative proabaility of -0.125

OpenStudy (ray_sihota):

so i look up -0.125 on the Z table

OpenStudy (reemii):

`What is the probability of getting a sample mean monthly rent that exceeds $660` Translation: \(\mathbb P\left( \bar X > 660 \right) = ?\) So, \(\mathbb P \left( \frac{(\bar X - 650)}{80/10} > \frac{(660 - 650)}{80/10} \right) = \mathbb P \left( \frac{(\bar X - 650)}{8} > \frac{10}{8} \right) = 1 - \mathbb P \left( \frac{(\bar X - 650)}{8} \le \frac{10}{8} \right) \) \(= \mathbb P( Normal(0,1) \le 1.25 = 1 - 0.8945\) That is the complete computation. You must learn to turn an inequality with \(\bar X\) into an inequality with \(\frac{\bar X - m}{\sigma/\sqrt{n}}\).

OpenStudy (reemii):

And use the cool behavior of \(\frac{\bar X - m}{\sigma/\sqrt n}\): it has a standard normal distribution. That's when you can use the table.

OpenStudy (ray_sihota):

question when do I use the T table

OpenStudy (reemii):

When the standard deviation is not given. Then, usually, the statement gives another value: an empirical measure of \(\sigma\), noted \(s\).

OpenStudy (reemii):

You must look carefully in your book or the statement. The "cool fraction" might show a \(n-1\) instead of an \(n\): \[\frac{\bar X-m}{s/\sqrt{n-1}}\]